# An overview of condensation phenomenon in deep learning

### v1

@@ 7.3 Reduction of network width
An approach to reduce a trained network can be readily proposed [ZZLX21]. If a neural network is in an extremely condensed state, neurons within the same layer that share the same output function can be replaced by a single equivalent neuron. This equivalent neuron would have the input weights of the original neurons and an output weight that is the sum of the output weights of the original neurons. Consequently, the original neural network can be reduced to a much narrower network, thereby saving computational costs during the inference stage. [CX24] utilize this reduction method for learning combustion problems, employing neural networks to solve ODEs through a data-driven approach. However, it should be noted that if a neural network is not in an extremely condensed state, such reduction can potentially harm performance, depending on the degree of condensation. Continuous training of the reduced network can mitigate this harm.

### v2

@@ 8.3 Reduction of network width
An approach to reduce a trained network can be readily proposed [ZZL+21]. If a neural network is in an extremely condensed state, neurons within the same layer that share the same output function can be replaced by a single equivalent neuron. This equivalent neuron would have the input weights of the original neurons and an output weight that is the sum of the output weights of the original neurons. Consequently, the original neural network can be reduced to a much narrower network, thereby saving computational costs during the inference stage. [CX24] utilize this reduction method for learning combustion problems, employing neural networks to solve ODEs through a data-driven approach. However, it should be noted that if a neural network is not in an extremely condensed state, such reduction can potentially harm performance, depending on the degree of condensation. Continuous training of the reduced network can mitigate this harm.